OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Circle Q has radius QR of length 55 cm. Point P is 73 cm. from Q, and PR is a tangent segment. How long is PR? a. 18 cm. b. 48 cm. c. 64 cm. d. 128 cm.

OpenStudy (anonymous):

QPR forms a right triangle. With PQ as the hypotenuse (PR)^2 = 73^2 - 55^2 (PR)^2 = 2304 Take the square root PR = 48